Hello I need help on this question, please explain how you got the the answer
"The sum of three consecutive odd integers is -87, what are the integers?"

Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

"The sum of three consecutive odd integers is -87, what are the integers?"
----------
(n-2) + n + (n+2) = -87

Answer by addingup(3677)

1) consecutive odd integers, example: 3,5,7 How many numbers in between them? the answer is 2
2) let the first integer be x. Then, the second will be x+2 and the third x+2+2 or simply x+4
Let's put it all together:
x+x+2+x+4 = -87
add like terms on left
3x+6 = -87
3x = -93
x = -31
------------------------
Check:
-31+-31+2+-31+4 = -87
-31+-29+-27 = -87
add the negatives and you'll see that you get -87 (some calculators don't like working with negative numbers, you can add them as positives and then add the negative sign in front of the sum).
:
John